Find the value of a+b+c if 1/a+1/b=c and a=b=c. Thanks for your help.
Answer by ewatrrr(24785)

1/a+1/b=c and a=b=c
.....
1/c + 1/c = c
1 + 1 = c^2
2 = c^2
c = ± √2 and a=b=c
....
a+b+c = ± 3√2